Question 1
BWS Corporation pays the premiums on an $80,000 group-term life insurance policy on the life of its 45-year-old vice-president, Warren. The annual cost per $1,000 of coverage for a person aged 45 to 49 is $1.80. If Warren has paid $25 toward the cost of the insurance, what is the cost of hte policy includible in Warren's gross income?

1) $278.00

2) $144.00

3) $54.00

4) $29.00

5) $0


Question 2
During 2019, Edward East had wages of $10,000 and received unemployment compensation of $6,200 from the state. Edward is single and 45 years old. What is the amount of unemployment compensation to be included in his gross income?

1) $0

2) $2,100

3) $3,800

4) $6,200


Question 3
Mr. and Mrs. Clark are both 72 years of age, married, filing jointly. They received social security benefits of $7,500 each. Additional income is a taxable pension of $12,000 for Mr. Clark and $6,000 for Mrs. Clark. What is the taxable portion of the social security benefits received by the Clarks?

1) $15,000

2) $7,500

3) $3,750

4) $0


Question 4
During the current year, Alfred Allen sustained a serious injury in the course of his employment. As a result of this injury, Allen received the following amounts during the same year:

Worker's compensation   $2,400
Reimbursement from employer's accident and health plan for medical expenses paid by Allen   1,800
Damages for personal injuries   8,000


How much of the above amounts should Allen include in his gross income for the current year?

1) $12,000

2) $8,000

3) $1,800

4) $0


Question 5
Amounts received under worker's compensation as compensation for personal injuries are excludable from gross income.
   1) True
   2) False


Question 6
In 2019, Bill, a single individual, earned a salary of $10,000 and received $2,000 in unemployment benefits. The unemployment benefits received by Bill are included in Bill's 2019 gross income.
   1) True
   2) False


Question 7
In 2019, Max is 85 years of age and single. He received Social Security payments totaling $14,000 (this includes Medicare premium of $600), dividend and interest income of $21,000, a pension of $30,000 and a taxable IRA benefits of $16,000. What is Max's adjusted gross income for 2019?

1) $80,400

2) $78,900

3) $78,390

4) $67,000


Question 8
Frank Clarke, an employee of Smithson Company, was covered under a noncontributory pension plan. Frank died on April 15, 2019, at age 64 and, pursuant to the plan, his widow received monthly pension payments of $500 beginning May 1, 2019. In addition, Mrs. Clarke received an employee death payment of $10,000 in May of 2019. This non-forfeitable death benefit was part of a group plan. What is the total amount of the above receipts that the widow should exclude from her gross income for 2019?

1) $0

2) $5,000

3) $9,000

4) $14,000


Question 9
Rachel Reeves has medical insurance coverage from her employer's medical insurance policy. The annual premium is $2,000 for Rachel's coverage, of which Rachel's employer pays 75 percent. Rachel pays the rest. In 2019, Rachel received an excess reimbursement of $200 from her insurance company. What amount must Rachel include in her gross income?

1) $0

2) $50

3) $75

4) $150

5) $200


Question 10
Dividend payments made by an insurance company that are based on a policy and that exceed the total amount of premiums paid by the insured are taxable to the insured.
   1) True
   2) False


Question 11
Mr. W. is 66 years old and single. His income for 2019 consisted of the following:


         
    Taxable pension   $10,000
    Taxable interest   2,000
    Taxable dividends   5,000
    Social security payments   5,000
    Tax-exempt interest   7,000


He did not have any adjustments to income. What amount of W's social security benefits is taxable?

1) $0

2) $750

3) $1,500

4) $2,000


Question 12
Claire is a sophomore in the University of Nebraska's degree program in dentistry. In 2019, Claire paid $3,000 in tuition, $500 for books, and $250 for rented dental equipment. Claire also paid room and board of $3,500. What is the total qualifying educational expense for Claire in 2019?

1) $3,500

2) $3,750

3) $7,250

4) $3,000


Question 13
Don Driller, who is 56 years old, is provided with $120,000 of group-term life insurance by his employer. Based on the IRS uniform premium cost table, the total annual cost of a policy of this type is $9.00 per $1,000 of coverage. Don's required contribution to the cost of the policy is $2.00 per $1,000 of coverage per year. Don was covered for the full 12 months of 2019. How much of the cost must Don include in his income for 2019?

1) $0

2) $390

3) $630

4) $840


Question 14
During 2019, Anne Apple received tangible personal property as a safety achievement award from her employer. The award was not a qualified plan award. The property cost the employer $500 and had a fair market value of $600. How much must Anne include in her 2019 gross income?

1) $0

2) $200

3) $500

4) $600
Question 15
During 2019, under a qualified written plan for educational assistance, Jake Jenner's employer paid Local University $9,200 on Jake's behalf: $7,500 for tuition, books, supplies, and lab fees and $1,700 for lodging. What is the amount of assistance that Jake should include in his gross income for 2019 if the payment was before May 1, 2019?

1) $9,200

2) $3,950

3) $2,350

4) $1,700
Answer #1

Dear Student,

As per the HOMEWORKLIB POLICY, only the first four questions should be answered. Kindly take note of it.

Part 1

Answer is option 4

4) $29.00

Total cost of policy = (80000-50000)/1000*1.80 = 54

Cost of policy includible in Warren's gross income = total cost of policy – cost paid = 54-25 $294

Part 2

Answer is option 4

4) $6,200

The full amount of unemployment compensation received will be included in the gross income

Part 3

Answer is option 2

2) $7,500

Total income = 15000+12000+6000 = $33000

It is less than threshold of $44000 and thus the taxable portion of social security benefits is 50%

Taxable portion = social security beneftis * 50% = 15000*50% = $7500

Part 4

Answer is option 4

4) $0

Benefits received such as worker’s compensation, compensation for damages for personal injuries as well as Reimbursement from employer's accident and health plan for medical expenses paid are excluded from gross income. For reimbursed expenses it is assumed that the medical expenses were not previously deducted as itemized deductions.